# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 3 Approximation of multipartite quantum states
$\Lambda^{s_{1},...,s_{l}}_{r}(\rho)=Q_{r}\rho\hskip 1.0ptQ_{r}[\mathrm{Tr}Q_{r}\rho\hskip 1.0pt]^{-1},\quad Q_{r}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{l}}\otimes I_{R},$ (24)
where $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account) and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$.
Theorem 1.  Let $A_{1}$,.,$A_{n}$ be arbitrary infinite-dimensional quantum systems, $n\geq 1$.
A) Let $\rho$ be a state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ such that $S(\rho_{A_{s}})<+\infty$ for $s=\overline{1,m}$, $m\leq n$. Let $\mathcal{F}\subset I_{n}^{m}$ be a family of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ having common faithful $I^{m}_{n}$-type continuity bound. [^11] See Section 2.3 (Definition 2.3 and below). Then for a given subset $\{s_{1},...,s_{l}\}$ of $\{1,...,n\}$ there is a vanishing sequence $\{Y(r)\}_{r=1}^{+\infty}$ such that
$|f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))-f(\rho)|\leq Y(r)$ (25)
for all $r\in\mathbb{N}$ and all functions $f\in\mathcal{F}$.
@@
$|f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))-f(\rho)|\leq Y(r)$ (26)
for all $r\in\mathbb{N}$ and all functions $f\in\mathcal{F}$.
The sequence $\{Y(r)\}_{r=1}^{+\infty}$ in claims A and B is completely determined by the states $\rho_{A_{s}}$, $s\in\{s_{1},...,s_{l}\}\cup\{1,...,m\}$, and the common continuity bound for the family $\mathcal{F}$.
Remark 3. In many applications of Theorem 3 the family $\mathcal{F}$ consists of a single function $f$. In this case it suffices to require that $f\in I_{n}^{m}$ in claim A and $f\in II_{n}^{m}$ in claim B.
Remark 4. Claim B of Theorem 3 differs from claim A by the weaker assumption about the family $\mathcal{F}$ (the class $II_{n}^{m}$ instead of $I_{n}^{m}$) which is compensated by the stronger assumption about the marginal states $\rho_{A_{s}}$, $s=\overline{1,m}$ (the FA-property instead of the finiteness of the entropy).
Claim B is more applicable to practical tasks, since the list of characteristics for which faithful $I_{n}^{m}$-type continuity bounds have been obtained (so far) is very short (see Section 2.3) while relation (18) (along with the observation in Section V in [31]) allows us to apply claim B to many characteristics of $n$-partite quantum systems.
Remark 5. The map $\Lambda^{s_{1},...,s_{l}}_{r}$ defined in (24) is nonlinear. The arguments from the proof of Theorem 3 show that its claims remain valid if we replace the family of maps $\Lambda^{s_{1},...,s_{l}}_{r}$ with the family of product-channels
@@
Proof of Theorem 3. A) For each natural $s$ in $[1,m]$ let $\{\varphi^{s}_{i}\}_{i=1}^{+\infty}$ be the basis of eigenvectors of the state $\rho_{A_{s}}$ corresponding to the non-increasing sequence $\{\lambda^{s}_{i}\}_{i=1}^{+\infty}$ of its eigenvalues. By Proposition 2.1 in Section 2.1 there is a positive operator
@@
on $\mathcal{H}_{A_{s}}$ satisfying condition (8) such that $g_{1}^{s}=0$ and
@@
$\|\rho-\Lambda^{s_{1},...,s_{l}}_{r}(\rho)\|_{1}\leq 2\sqrt{\mathrm{Tr}\bar{Q}_{r}\rho}\leq 2\sqrt{\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}},$ (31)
@@
Let $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ be a common faithful $I_{n}^{m}$-type continuity bound for the family $\mathcal{F}$. Let $f$ be any function in $\mathcal{F}$. Since $\,\sum_{s=1}^{m}\mathrm{Tr}G_{A_{s}}\rho_{A_{s}}=E_{S}$ and all the operators $G_{A_{1}}$,…,$G_{A_{m}}$ satisfy condition (8), by taking into account inequalities (30) and (31) we obtain
$|f(\rho)-f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))|\leq\widehat{B}_{\mathcal{F}}(\varepsilon_{r},2E_{S}|G_{\!A_{1}},..,G_{\!A_{m}}),\quad\varepsilon_{r}=\sqrt{\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}},$ (32)
for all $r>r_{0}$, where $r_{0}$ is some natural number.
Since $\varepsilon_{r}\to+\infty$ as $r\to+\infty$, the r.h.s. of (32) tends to zero as $r\to+\infty$. By denoting the r.h.s. of (32) by $Y(r)$ for $r>r_{0}$ and assuming that
$Y(r)=\sup_{f\in\mathcal{F}}|f(\rho)-f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))|$ (33)
for $r\leq r_{0}$ we obtain the claim A of the theorem (by using the common continuity bound $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ and inequality (30) it is easy to show that the supremum in (33) is finite for any $r$).
B) For each natural $s$ in $[1,m]$ let $\{g^{s}_{i}\}_{i=1}^{+\infty}$ be a sequence of nonnegative numbers such that
$\sum_{i=1}^{+\infty}\lambda^{s}_{i}g^{s}_{i}<+\infty\quad\textrm{and}\quad\lim_{\beta\to 0^{+}}\left[\sum_{i=1}^{+\infty}e^{-\beta g^{s}_{i}}\right]^{\beta}=1,$ (34)
where $\{\lambda^{s}_{i}\}_{i=1}^{+\infty}$ is the spectrum of the state $\rho_{A_{s}}$ taking in the non-increasing order. We may assume that $g_{1}^{s}=0$ for all $s$.
Then, by defining for each $s$ the positive operator $G_{\!A_{s}}$ on $\mathcal{H}_{A_{s}}$ by the formula (28) we may repeat all the arguments of the proof of claim A with the use of the common faithful $II_{n}^{m}$-type continuity bound $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for the family $\mathcal{F}$ at the last step (since now all the operators $G_{A_{1}}$,…,$G_{A_{m}}$ satisfy condition (10)).
By noting that the r.h.s. of (32) of tends to zero as $r\to+\infty$ and by constructing the sequence $Y(r)$ in the same way as in the proof of claim A we obtain claim B of the theorem.
@@
Example 3. Let $f_{\Phi_{1}...\Phi_{n}}(\rho)=I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}$, where $\Phi_{1}:A_{1}\rightarrow B_{1}$,…, $\Phi_{n}:A_{n}\rightarrow B_{n}$ are arbitrary quantum channels. Inequality (10) in [31] (with trivial system $C$), upper bound (4), the nonnegativity of the quantum mutual information and its monotonicity under local channels imply that the function $f_{\Phi_{1}...\Phi_{n}}$ belongs to the class $L_{n}^{n-1}(2,n)$ for any channels $\Phi_{1},...,\Phi_{n}$. [^13] See the remark at the end of the Appendix. It follows that the quantity $\widehat{B}_{\mathcal{F}(2,n)}(\varepsilon,E|H_{A_{1}},..,H_{A_{n-1}})$ defined in (19) is a common faithful $II_{n}^{n-1}$-type continuity bound for the family =$\mathcal{F}=\{f_{\Phi_{1}...\Phi_{n}}\}_{\Phi_{1}...\Phi_{n}}$ (where $\Phi_{k}$ runs over the set of all channels from $A_{k}$ to $B_{k}$, $k=\overline{1,n}$). Thus, if the marginal states $\rho_{A_{1}}$,…,$\rho_{A_{n-1}}$ have the FA-property then, by Theorem 3B, for any subset $\{s_{1},...,s_{l}\}$ of $\{1,...,n\}$ there exists a vanishing sequence $\{Y(r)\}$ such that
$|I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho_{r})}-I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}|\leq Y(r)\qquad\forall r\in\mathbb{N},$ (35)
@@
The last assertion of Theorem 3 shows that inequality (35) remains valid with $\rho$ and $\rho_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\rho)$ replaced by $\sigma$ and $\sigma_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\sigma)$, where $\sigma$ is any state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ such that $\sigma_{A_{k}}=\rho_{A_{k}}$ for each $k\in\{s_{1},...,s_{l}\}\cup\{1,...,n-1\}$.
By using expression (23), the Stinespring representation of the channels $\Phi_{1}$,…,$\Phi_{n-1}$ and Proposition 7 in [33, Section 4.2] it is easy to show the uniform continuity of the function
$(\Phi_{1},...,\Phi_{n})\mapsto I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho_{r})}$
on the set of all tuples $(\Phi_{1},...,\Phi_{n})$ of channels equipped with the topology of pairwise strong convergence, where $\rho_{r}=\Lambda^{1,...,n-1}_{r}(\rho)$, for any $r$. Thus, the above observation with $\{s_{1},...,s_{l}\}=\{1,...,n-1\}$ implies continuity of the function
$(\Phi_{1},...,\Phi_{n})\mapsto I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}$
on the set of all tuples $(\Phi_{1},...,\Phi_{n})$ of channels equipped with the topology of pairwise strong convergence provided that the FA-property holds for at least $n-1$ marginal states of the state $\rho$. By noting that the FA-property implies the finiteness of the entropy, this assertion can be also obtained from Proposition 3 in [30] proved in a completely different way.
4 On continuity of the regularized relative entropy of $\pi$-entanglement
In this section we apply Theorem 1 along with the results from [15] to obtain conditions of local continuity for the regularized relative entropy of $\pi$-entanglement of a state of $n$-partite system $A_{1}\ldots A_{n}$ corresponding to any given partition $\pi$ of $[n]\doteq\{1,\ldots,n\}$ (i.e. a finite collection $\{\pi(j)\}_{j\in J}$ of non-empty sets that do not intersect each other and together cover the whole set $[n]$). For a given $\pi$ we define the set of $\pi$-separable states by (cf.[25])

### v2

@@ 3.1 Basic theorem
$\Lambda^{s_{1},...,s_{l}}_{r}(\rho)=[\mathrm{Tr}Q_{r}\rho\hskip 1.0pt]^{-1}Q_{r}\rho\hskip 1.0ptQ_{r},\quad Q_{r}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{l}}\otimes I_{R},$ (27)
where $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account) and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$. [^11] Here and in what follows speaking about the spectral projector $P_{r}$ of a state $\rho$ corresponding to its $r$ maximal eigenvalues we assume that $P_{r}$ is the projector onto the subspace $\mathrm{supp}\rho$ if $\,r>\mathrm{rank}\rho$.
Theorem 1.  Let $A_{1}$,..,$A_{n}$ be arbitrary infinite-dimensional quantum systems, $n\geq 1$
A) If $f$ is a function on $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ having faithful $I^{m}_{n}$-type continuity bound $\widehat{B}_{f}$ [^12] See Section 2.3 (Definition 2.3 and below). then for any state $\rho$ in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ such that $S(\rho_{A_{s}})<+\infty$ for $s=\overline{1,m}$ the following properties hold:
  • $f(\rho)$ is finite;
  • for an arbitrary subset $\{s_{1},...,s_{l}\}$ of $\,\{1,...,n\}$ there is a sequence $\{Y_{r}\}_{r=1}^{+\infty}$ tending to zero such that
    $|f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))-f(\rho)|\leq Y_{r}\quad\forall r\in\mathbb{N}.$ (28)
@@
The sequence $\{Y_{r}\}_{r=1}^{+\infty}$ in claims A and B is completely determined by the states $\rho_{A_{s}}$, $s\in\{s_{1},...,s_{l}\}\cup\{1,...,m\}$, and by the continuity bound $\widehat{B}_{f}$.
Remark 3. Claim B of Theorem 3.1 differs from claim A by the weaker assumption about the function $f$ (the class $II_{n}^{m}$ instead of $I_{n}^{m}$) which is compensated by the stronger assumption about the marginal states $\rho_{A_{s}}$, $s=\overline{1,m}$ (the FA-property instead of the finiteness of the entropy).
Claim B is more applicable to practical tasks, since the list of characteristics for which faithful $I_{n}^{m}$-type continuity bounds have been obtained (so far) is quite short, while relation (24) (along with the observation in Section V in [15]) allows us to apply claim B to many characteristics of $n$-partite quantum systems (non-complete lists of characteristics belonging to the classes $I_{n}^{m}$ and $II_{n}^{m}$ are placed in the Appendix). Note that claim B can be applied to all characteristics from the class $I_{n}^{m}$ as well, since $I_{n}^{m}\subset II_{n}^{m}$.
Remark 4. The map $\Lambda^{s_{1},...,s_{l}}_{r}$ defined in (27) is nonlinear. The arguments from the below proof of Theorem 3.1 show that its claims remain valid if we replace the family of maps $\Lambda^{s_{1},...,s_{l}}_{r}$ with the family of product-channels
@@
Proof of Theorem 3.1. A) For each natural $s$ in $[1,m]$ let $\mathcal{T}_{s}\doteq\{\varphi^{s}_{i}\}_{i=1}^{+\infty}$ be an orthonormal system of eigenvectors of the state $\rho_{A_{s}}$ corresponding to the non-increasing sequence $\{\lambda^{s}_{i}\}_{i=1}^{+\infty}$ of its eigenvalues (if $r\doteq\mathrm{rank}\rho_{A_{s}}<+\infty$ then $\mathcal{T}_{s}$ is an arbitrary extension of the basis $\{\varphi^{s}_{i}\}_{i=1}^{r}$ of eigenvectors of $\rho_{A_{s}}$ in $\,\mathrm{supp}\rho_{A_{s}}$ to a countable orthonormal system of vectors and $\lambda^{s}_{i}=0$ for all $i>r$). By Proposition 2.1 in Section 2.1 (and its proof) there is a positive operator
@@
on $\mathcal{D}(G_{\!A_{s}})=\{\psi\in\mathrm{supp}\rho_{A_{s}}\,|\,\sum_{i=1}^{+\infty}g^{2}_{i}|\langle\varphi^{s}_{i}|\psi\rangle|^{2}<+\infty\}\subseteq\mathcal{H}_{A_{s}}$ satisfying condition (15) such that $g_{1}^{s}=0$ and
@@
$\|\rho-\Lambda^{s_{1},...,s_{l}}_{r}(\rho)\|_{1}\leq 2\sqrt{\mathrm{Tr}\bar{Q}_{r}\rho}\leq 2\sqrt{\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}}=o(1)\quad\textrm{ as }\;r\to+\infty,$ (33)
@@
Let $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ be a faithful $I_{n}^{m}$-type continuity bound for the function $f$. Since $\,\sum_{s=1}^{m}\mathrm{Tr}G_{A_{s}}\rho_{A_{s}}=E_{S}$ and all the operators $G_{A_{1}}$,…,$G_{A_{m}}$ satisfy condition (15), we conclude that $f(\rho)$ is finite and, by taking into account (32) and (33), we obtain
$|f(\rho)-f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))|\leq\widehat{B}_{f}(\varepsilon_{r},2E_{S}|\hskip 1.0ptG_{\!A_{1}},..,G_{\!A_{m}}),\quad\varepsilon_{r}=\sqrt{\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}},$ (34)
for all $r>r_{0}$, where $r_{0}$ is some natural number (not depending on $f$).
Since $\varepsilon_{r}\to 0$ as $r\to+\infty$, the r.h.s. of (34) tends to zero as $r\to+\infty$. By denoting the r.h.s. of (34) by $Y_{r}$ for $r>r_{0}$ and setting $Y_{r}=\widehat{B}_{f}(\varepsilon_{r},E_{r}|\hskip 1.0ptG_{\!A_{1}},..,G_{\!A_{m}})$ for $r\leq r_{0}$, where $E_{r}=E_{S}/(1-\varepsilon^{2}_{r})$, we obtain claim A of the theorem.
B) By the definition of the FA-property for each natural $s$ in $[1,m]$ there exists a sequence $\{g^{s}_{i}\}_{i=1}^{+\infty}$ of nonnegative numbers such that
$\sum_{i=1}^{+\infty}\lambda^{s}_{i}g^{s}_{i}<+\infty\quad\textrm{and}\quad\lim_{\beta\to 0^{+}}\left[\sum_{i=1}^{+\infty}e^{-\beta g^{s}_{i}}\right]^{\beta}=1,$
where $\{\lambda^{s}_{i}\}_{i=1}^{+\infty}$ is the sequence of eigenvectors of the state $\rho_{A_{s}}$ arranged in the non-increasing order. We may assume that $g_{1}^{s}=0$ for all $s$.
Then, by defining for each $s$ the positive operator $G_{\!A_{s}}$ on $\mathcal{H}_{A_{s}}$ by the formula (30) we may repeat all the arguments of the proof of claim A with the use of the faithful $II_{n}^{m}$-type continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for the function $f$ at the last step (since now all the operators $G_{A_{1}}$,…,$G_{A_{m}}$ satisfy condition (17)).
By noting that the r.h.s. of (34) tends to zero as $r\to+\infty$ and by constructing the sequence $Y_{r}$ in the same way as in the proof of claim A we obtain claim B of the theorem.
@@
Remark 5. If $\mathcal{F}$ is a family of functions having a common faithful $I^{m}_{n}$-type (resp. $II^{m}_{n}$-type) continuity bound then the last claim of Theorem 3.1 implies that inequality (28) holds for all functions $f\in\mathcal{F}$ with the sequence $\{Y_{r}\}_{r=1}^{+\infty}$ tending to zero and not depending on $f$. So, in this case the approximation given by Theorem 3.1 is uniform.
Example 3. Let $f_{\Phi_{1}...\Phi_{n}}(\rho)=I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}$, where $\Phi_{1}:A_{1}\rightarrow B_{1}$,…, $\Phi_{n}:A_{n}\rightarrow B_{n}$ are arbitrary quantum channels. The arguments from Section V in [15] show the function $f_{\Phi_{1}...\Phi_{n}}$ belongs to the class $L_{n}^{n-1}(2,n)$ for any channels $\Phi_{1},...,\Phi_{n}$. It follows that the quantity $\widehat{B}_{2,n}(\varepsilon,E|H_{A_{1}},..,H_{A_{n-1}})$ defined in (25) is a common faithful $II_{n}^{n-1}$-type continuity bound for the family =$\mathcal{F}=\{f_{\Phi_{1}...\Phi_{n}}\}_{\Phi_{1}...\Phi_{n}}$ (where $\Phi_{k}$ runs over the set of all channels from $A_{k}$ to $B_{k}$, $k=\overline{1,n}$). Thus, if the marginal states $\rho_{A_{1}}$,…,$\rho_{A_{n-1}}$ have the FA-property then Theorem 3.1B and Remark 3.1 imply that for any subset $\{s_{1},...,s_{l}\}$ of $\{1,...,n\}$ there exists a sequence $\{Y_{r}\}$ tending to zero such that
$|I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho_{r})}-I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}|\leq Y_{r}\qquad\forall r\in\mathbb{N},$ (35)
@@
The last assertion of Theorem 3.1 shows that inequality (35) remains valid with $\rho$ and $\rho_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\rho)$ replaced by $\sigma$ and $\sigma_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\sigma)$, where $\sigma$ is any state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ such that $\sigma_{A_{k}}=\rho_{A_{k}}$ for each $k\in\{s_{1},...,s_{l}\}\cup\{1,...,n-1\}$.
3.2 Some corollaries
By using the results of the previous subsection one can obtain the following observations that can be applied to analysis of entanglement and correlations measures in multipartite quantum systems (concrete applications are considered in Section 4).
Corollary 1. Let $E$ be a nonnegative function on the set $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ such that
$E(\Phi_{1}\otimes...\otimes\Phi_{m}\otimes\mathrm{Id}_{A_{m+1}...A_{n}}(\rho))\leq E(\rho)$ (36)
for any $\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and any quantum channels $\Phi_{s}:A_{s}\to A_{s}$, $s=\overline{1,m}$, $m\leq n$. Let $\,\mathfrak{S}_{0}(\mathcal{H}_{A_{1}...A_{n}})$ be the trace norm closure of the set $E^{-1}(0)\cap\mathfrak{S}_{\rm f}(\mathcal{H}_{A_{1}...A_{n}})$, where